Find the quotient field of integral domain {a+ib such that a,b belongs to Z}
1
Expert's answer
2020-10-14T18:16:44-0400

Let "K" be the quotient field of integral domain "\mathbb Z[i]=\{a+bi\ |\ a,b\in\mathbb Z\}" and let "\mathbb Q(i)" the field of complex numbers of the form "r+si" with both "r" and "s" in "\mathbb Q" . Let us prove that "K=\mathbb Q(i)".


First thing we need to convince yourseld that "K\subset\mathbb Q(i)". If we take a fraction "\frac{a+bi}{c+di}" with "a,b,c,d\in\mathbb Z" we can rewrite it as "\frac{a+bi}{c+di}=\frac{ac+bd}{c^2+d^2}+\frac{bc-ad}{c^2+d^2}i", and we're done.


On the other hand, if "\frac{m}{n}+\frac{p}{q}i\in\mathbb Q(i)", we can rewrite this as "\frac{nq+npi}{nq}" where both numerator and denominator are Gaussian integers. Thus we also have "\mathbb Q(i)\subset K" and we're finished.